# references.bib

### v2

@@ Abstract
Multi-objective Bayesian optimization (MOBO) provides a principled framework for navigating trade-offs in molecular design. However, its empirical advantages over scalarized alternatives remain underexplored. We benchmark a simple Pareto-based MOBO strategy—Expected Hypervolume Improvement (EHVI)—against a simple fixed-weight scalarized baseline using Expected Improvement (EI), under a tightly controlled setup with identical Gaussian Process surrogates and molecular representations. Across three molecular optimization tasks, EHVI consistently outperforms scalarized EI in terms of Pareto front coverage, convergence speed, and chemical diversity. While scalarization encompasses flexible variants—including random or adaptive schemes—our results show that even strong deterministic instantiations can underperform in low-data regimes. These findings offer concrete evidence for the practical advantages of Pareto-aware acquisition in de novo molecular optimization, especially when evaluation budgets are limited and trade-offs are nontrivial.
@@ 1 Introduction
The discovery of therapeutic molecules is fundamentally a multi-objective optimization problem: a viable candidate must simultaneously satisfy competing criteria such as potency, safety, and pharmacokinetic properties [nicolaou2013multi, ekins2010evolving, fromer2023computer]. Scalarization—where multiple objectives are collapsed into a single score using weighted combinations [murata1996multiscalarization, helfrich2024scalarization] - remains a widely used strategy due to its compatibility with single-objective optimization pipelines. However, scalarization requires a priori knowledge of objective weightings, which are often uncertain, context-dependent, or poorly defined in real-world drug design. Furthermore, any fixed weighting yields only a single point on the Pareto front, necessitating repeated and often redundant optimization runs to recover a diverse set of trade-off solutions.
These limitations motivate Pareto-based multi-objective Bayesian optimization (MOBO) methods that preserve vector-valued structure of the problem and directly seek non-dominated solutions across all objectives[fromercoleypareto2024, zhu2023sample, ahmadbelakariamoo2024]. Rather than collapsing objectives, MOBO aims to efficiently approximate the Pareto front by guiding evaluations toward regions that improve coverage. This approach has been shown to recover more chemically diverse and balanced solutions using fewer queries - particularly valuable in low-data, expensive-to-evaluate regimes. Beyond molecular design, MOBO has also demonstrated success across domains including materials science [gopakumar2018multi, kumar2022multiobjective], protein engineering [park2022propertydag], and robotics [kim2021mo, wang2024fin].
Despite this growing interest, few empirical studies systemically benchmark Pareto-based MOBO against specific scalarization strategies under controlled conditions. Prior work often defaults to scalarization heuristics [fromer2023computer, jain2023multi] without evaluating their performance relative to the dedicated Pareto-based acquisitions. In this study, we present a controlled comparison between a fixed-weight scalarized Bayesian optimization strategy (using Expected Improvement) and a Pareto-based MOBO approach (using Expected Hypervolume Imrprovement). Both are implemented using identical Gaussian Process surrogates and molecular representations, isolating the acquisition function as the key difference. By evaluating performance across three GUACAMOL benchmark tasks, we highlight how even basic Pareto-based strategies can outperform scalarization in data-constrained molecular discovery scenarios - supporting MOBO as a more robust default for early-stage optimization.
@@ 2.2 Bayesian Optimization with Gaussian Processes
Bayesian optimization is a sample-efficient framework for optimizing expensive black-box functions. It employs a surrogate model - typically a Gaussian Process ($\mathcal{GP}$) - to model the function’s uncertainty and selects new evaluations using an acquisition function $\alpha(x)$ that balances exploration and exploitation [brochu2010].
A Gaussian Process ($\mathcal{GP}$) is the most widely used surrogate in BO due to its nonparametric flexibility and closed-form posterior. A $\mathcal{GP}$ prior over functions $f\colon\mathcal{X}\to\mathbb{R}$ is specified by a mean function $m(\mathbf{x})$ (often taken as zero) and a covariance (kernel) function $k(\mathbf{x},\mathbf{x}^{\prime})$ [rassmussen2006]. Given observations $\mathcal{D}_{n}={(\mathbf{x}i,y_{i})}{i=1}^{n}$ with Gaussian noise $y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i},;\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, the $\mathcal{GP}$ posterior at a candidate $\mathbf{x}$ has predictive distribution:
@@
with $K_{n}[i,j]=k(\mathbf{x}_{i},\mathbf{x}_{j})$ and $\mathbf{k}_{n}(\mathbf{x})=[k(\mathbf{x},\mathbf{x}i)]{i=1}^{n}$ [rassmussen2006]. These expressions quantify both the surrogate mean and its epistemic uncertainty, enabling a principled trade-off in $\alpha(\mathbf{x})$.
@@
$k_{Tanimoto}(x,x^{\prime})=\frac{x^{\top}x^{\prime}}{||x||_{2}^{2}+||x^{\prime}||^{2}_{2}-x^{\top}x^{\prime}}$
@@ 2.3 Scalarization and Pareto-based Optimization Strategies
To overcome these limitations, alternative scalarization techniques such as the Tchebycheff scalarization have been developed $f_{Tchebycheff}(x)=\max_{1\leq i\leq m}\{\lambda_{i}(f_{i}(x)=z^{*}_{i})\}$ where $\lambda_{i}$ are scaling weights and $z_{i}^{*}$ is a reference point (typically the ideal vector of component-wise maxima). This scalarization emphasizes the worst-performing objective, yielding stronger guarantees for recovering solutions on convex fronts [tchebysheff1976, smoothsettchebysheff2025]. Recent work has shown that only a small set of well-chosen Tchebycheff scalarizations can approximate the entire Pareto front with high fidelity and sample efficiency [smoothsettchebysheff2025].
A particularly impactful advancement comes from the hypervolume scalarization framework introduced by Golovin and Zhang (2020)[golovin2020randomhypervolumescalarizationsprovable], which establishes a formal connection between scalarized objectives and the hypervolume indicator — a gold-standard, Pareto-compliant metric [zitzler1998multiobjective]. Specifically, they show that for a suitable distribution $\mathcal{D}_{\lambda}$ over weight vectors and corresponding scalarization functions $s_{\lambda}$, the hypervolume of a set $Y\subset\mathbb{R}^{d}$ with respect to a reference point $z\in\mathbb{R}^{d}$ can be expressed as $HV_{z}(Y)=c_{k}\,\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\max_{y\in Y}s_{\lambda}(y-z)\right]$.
This theoretical result provides a principled mechanism for converting hypervolume maximization into a sequence of scalar optimization subproblems. In practice, it enables provable convergence to the Pareto front using standard Bayesian optimization techniques, such as Thompson Sampling or UCB, by simply sampling a new scalarization $s_{\lambda}$ at each iteration. Golovin and Zhang (2020) derive cumulative hypervolume regret bounds of order $\tilde{O}(T^{-1/2})$, establishing the method as both theoretically grounded and sample-efficient for multi-objective black-box optimization.
@@ 3 Experimental Setup
Our experiments are designed to assess the practical performance of Pareto-based acquisition strategies for multi-objective Bayesian optimization (MOBO) in molecular design tasks. Rather than contrasting MOBO with scalarization broadly—which encompasses a diverse array of formulations including random scalarizations [golovin2020randomhypervolumescalarizationsprovable, paria2019flexibleframeworkmultiobjectivebayesian] - we focus on a controlled comparison between two representative acquisition strategies: Expected Hypervolume Improvement (EHVI), which explicitly targets Pareto front expansion, and a fixed-weight Expected Improvement (EI), a baseline scalarized acquisition function. We aim to answer the following questions:
@@ 3.3 Evaluation Metrics and Performance Indicators
To evaluate the performance of multi-objective optimization across both convergence quality and molecular diversity, we adopted three complementary metrics: Hypervolume Indicator (HVI)[fonsecahypervolume], the R2 indicator [hansenr2] and the #Circles metric [xie2023spaceexploredmeasuringchemical]. The HVI measures the volume of the objective space dominated by the non-dominated solutions relative to a reference point, capturing both convergence to the Pareto front and diversity of trade-offs explored. A higher HVI value indicates broader and more optimal coverage of the objective space. The $R^{2}$ indicator quantifies the quality of the Pareto front approximation by comparing it to a fixed set of uniformly distributed reference directions, as adapted from Jain et al(2023)’s[jain2023multi] setup. Specifically, it computes the augmented Chebyshev scalarization in each direction: for each reference vector $v$, it calculates the worst-case deviation $\text{max}_{i}v_{i}\cdot|u_{i}-s_{i}|$, where $u$ is the utopian (ideal) point and $s$ is a candidate solution. It then selects the best such solution for each direction and averages over all directions. Thus, lower $R^{2}$ values indicate a front that is both closer and more uniformly distributed relative to the ideal front [hansenr2]. Finally, the #Circles metric quantifies structural diversity in the chemical space by counting the number of pairwise dissimilar molecules exceeding a Tanimoto distance threshold $t$. We compute it on the set of Pareto-optimal candidates from the initial and acquired molecules over 200 BO evaluations, as the metric is designed to assess the effective space explored by a representative set of high-quality or relevant solutions rather than all generated samples [xie2023spaceexploredmeasuringchemical]. Informed by axiomatic diversity principles, #Circles offers a geometry-aware, thresholded view of chemical diversity: at higher thresholds (i.e. requiring greater dissimilarity), a larger #Circles value indicates that the candidate set spans more structurally distinct regions of the chemical space.
@@ 4 Results
This section presents a comparative evaluation of EHVI and scalarized EI across 3 multi-objective molecular optimization tasks. Across all tasks and metrics - hypervolume (Figure 1), $R^{2}$ indicator (Figure 2), and chemical diversity via the #Circles metric (Figure 3), EHVI demonstrates consistent advantages. Statistical tests provided in Appendix 7.2 (Tables 3, 4) corroborate these trends: for hypervolume, EHVI achieves medium to large effect sizes (Cohen’s $d=0.576-1.093$) and favorable Cliff’s delta values indicating better performance across matched trials. For the $R^{2}$ indicator, lower values for EHVI suggest improved approximation of the Pareto front, with strong negative effect sizes (e.g. $d=-2.56$ on Fexofenadine MPO). In terms of diversity, EHVI consistently explores more structurally distinct solutions at higher Tanimoto thresholds. These results demonstrate EHVI’s robustness: it not only accelerates convergence and enhances front coverage, it also maintains better front approximation and promotes chemically diverse candidate solutions. We next present the detailed metric-wise breakdown across all tasks.
@@ 5 Discussion
Our findings demonstrate that EHVI yields more sample-efficient exploration across a range of realistic multi-property optimization (MPO) tasks. These gains are reflected across three orthogonal metrics—hypervolume coverage, front approximation accuracy ($R^{2}$), and structural diversity—and are statistically robust across random seeds, as shown in Appendix 7.2 (Tables 3, 4). Crucially, these improvements are not due to changes in surrogate fidelity, kernel choice, or representation capacity: both EHVI and scalarized EI operate under the same model assumptions and input features, underscoring the importance of acquisition strategy alone in driving performance.
@@ 6 Conclusion
We conducted a focused comparison between Pareto-based Bayesian optimization (EHVI) and scalarized Expected Improvement (EI) with fixed weights across multi-objective molecular design tasks. EHVI consistently outperforms EI, achieving higher hypervolume, lower $R^{2}$ values, and greater structural diversity. Statistical effect size analysis further confirms these gains are meaningful despite limited runs. Unlike scalarized EI, which optimizes a fixed utility, EHVI explicitly captures trade-offs, leading to more balanced and diverse Pareto fronts. These results demonstrate the value of Pareto-aware acquisition in data-constrained settings and motivate future work on adaptive surrogates, representation learning, and diversity-aware BO strategies. Code repository is available here at https://github.com/anabelyong/efficient-mobo.
[…]
7 Appendix
7.1 Final Performance Metrics after 200 BO evaluations
@@
7.2 Cohen’s $d$ and Cliff’s Delta
@@
Fexofenadine $-2.560$ $-1.000$
Amlodipine $-0.770$ $-0.556$
Perindopril $-1.602$ $-0.778$

### v3

# A study of EHVI vs fixed scalarization for molecule design

@@ Abstract
Multi-objective Bayesian optimization (MOBO) provides a principled framework for navigating trade-offs in molecular design. However, its empirical advantages over scalarized alternatives remain underexplored. We benchmark a simple Pareto-based MOBO strategy—Expected Hypervolume Improvement (EHVI)-against a simple fixed-weight scalarized baseline using Expected Improvement (EI), under a tightly controlled setup with identical Gaussian Process surrogates and molecular representations. Across three molecular optimization tasks, EHVI consistently outperforms scalarized EI in terms of Pareto front coverage, convergence speed, and chemical diversity. While scalarization encompasses flexible variants - including random or adaptive schemes—our results show that even strong deterministic instantiations can underperform in low-data regimes. These findings offer concrete evidence for the practical advantages of Pareto-aware acquisition in de novo molecular optimization, especially when evaluation budgets are limited and trade-offs are nontrivial.
@@ 1 Introduction
The discovery of therapeutic molecules is fundamentally a multi-objective optimization problem: a viable candidate must simultaneously satisfy competing criteria such as potency, safety, and pharmacokinetic properties [nicolaou2013multi, ekins2010evolving, fromer2023computer]. Scalarization—where multiple objectives are collapsed into a single score using weighted combinations [murata1996multiscalarization, helfrich2024scalarization]—remains a widely used strategy due to its compatibility with single-objective optimization pipelines. However, scalarization requires a priori knowledge of objective weightings, which are often uncertain, context-dependent, or poorly defined in real-world drug design. Furthermore, any fixed weighting yields only a single point on the Pareto front, necessitating repeated and often redundant optimization runs to recover a diverse set of trade-off solutions.
These limitations motivate Pareto-based multi-objective Bayesian optimization (MOBO) methods that preserve vector-valued structure of the problem and directly seek non-dominated solutions across all objectives [fromercoleypareto2024, zhu2023sample, ahmadbelakariamoo2024]. Rather than collapsing objectives, MOBO aims to efficiently approximate the Pareto front by guiding evaluations toward regions that improve coverage. This approach has been shown to recover more chemically diverse and balanced solutions using fewer queries—particularly valuable in low-data, expensive-to-evaluate regimes. Beyond molecular design, MOBO has also demonstrated success across domains including materials science [gopakumar2018multi, kumar2022multiobjective], protein engineering [park2022propertydag], and robotics [kim2021mo, wang2024fin].
Despite this growing interest, few empirical studies systemically benchmark Pareto-based MOBO against specific scalarization strategies under controlled conditions. Prior work often defaults to scalarization heuristics [fromer2023computer, jain2023multi] without evaluating their performance relative to the dedicated Pareto-based acquisitions. In this study, we present a controlled comparison between a fixed-weight scalarized Bayesian optimization strategy (using Expected Improvement) and a Pareto-based MOBO approach (using Expected Hypervolume Imrprovement). Both are implemented using identical Gaussian Process surrogates and molecular representations, isolating the acquisition function as the key difference. By evaluating performance across three GUACAMOL benchmark tasks, we highlight how even basic Pareto-based strategies can outperform scalarization in data-constrained molecular discovery scenarios—supporting MOBO as a more robust default for early-stage optimization. To ensure reproducibility, we will release the code and data upon acceptance, with links provided in the final version of the paper.
@@ 2.2 Bayesian Optimization with Gaussian Processes
Bayesian optimization is a sample-efficient framework for optimizing expensive black-box functions. It employs a surrogate model—typically a Gaussian Process ($\mathcal{GP}$)—to model the function’s uncertainty and selects new evaluations using an acquisition function $\alpha(x)$ that balances exploration and exploitation [brochu2010].
A Gaussian Process ($\mathcal{GP}$) is the most widely used surrogate in BO due to its nonparametric flexibility and closed-form posterior. A $\mathcal{GP}$ prior over functions $f\colon\mathcal{X}\to\mathbb{R}$ is specified by a mean function $m(\mathbf{x})$ (often taken as zero) and a covariance (kernel) function $k(\mathbf{x},\mathbf{x}^{\prime})$ [rassmussen2006]. Given observations $\mathcal{D}_{n}={(\mathbf{x}_{i},y_{i})}_{i=1}^{n}$ with Gaussian noise $y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i},;\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, the $\mathcal{GP}$ posterior at a candidate $\mathbf{x}$ has predictive distribution:
@@
with $K_{n}[i,j]=k(\mathbf{x}_{i},\mathbf{x}_{j})$ and $\mathbf{k}_{n}(\mathbf{x})=[k(\mathbf{x},\mathbf{x}_{i})]_{i=1}^{n}$ [rassmussen2006]. These expressions quantify both the surrogate mean and its epistemic uncertainty, enabling a principled trade-off in $\alpha(\mathbf{x})$.
@@
[…]
@@ 2.3 Scalarization and Pareto-based Optimization Strategies
To overcome these limitations, alternative scalarization techniques such as the Tchebycheff scalarization have been developed, with $f_{Tchebycheff}(x)=\max_{1\leq i\leq m}\{\lambda_{i}(f_{i}(x)=z^{*}_{i})\}$ where $\lambda_{i}$ are scaling weights and $z_{i}^{*}$ is a reference point (typically the ideal vector of component-wise maxima). This scalarization emphasizes the worst-performing objective, yielding stronger guarantees for recovering solutions on convex fronts [tchebysheff1976, smoothsettchebysheff2025]. Recent work has shown that only a small set of well-chosen Tchebycheff scalarizations can approximate the entire Pareto front with high fidelity and sample efficiency [smoothsettchebysheff2025].
A particularly impactful advancement comes from the hypervolume scalarization framework introduced by golovin2020randomhypervolumescalarizationsprovable, which establishes a formal connection between scalarized objectives and the hypervolume indicator—a gold-standard, Pareto-compliant metric [zitzler1998multiobjective]. Specifically, they show that for a suitable distribution $\mathcal{D}_{\lambda}$ over weight vectors and corresponding scalarization functions $s_{\lambda}$, the hypervolume of a set $Y\subset\mathbb{R}^{d}$ with respect to a reference point $z\in\mathbb{R}^{d}$ can be expressed as $HV_{z}(Y)=c_{k}\,\mathbb{E}_{\lambda\sim\mathcal{D}_{\lambda}}\left[\max_{y\in Y}s_{\lambda}(y-z)\right]$.
This theoretical result provides a principled mechanism for converting hypervolume maximization into a sequence of scalar optimization subproblems. In practice, it enables provable convergence to the Pareto front using standard Bayesian optimization techniques, such as Thompson Sampling or UCB, by simply sampling a new scalarization $s_{\lambda}$ at each iteration. golovin2020randomhypervolumescalarizationsprovable derive cumulative hypervolume regret bounds of order $\tilde{O}(T^{-1/2})$, establishing the method as both theoretically grounded and sample-efficient for multi-objective black-box optimization.
@@ 3 Experimental Setup
Our experiments are designed to assess the practical performance of Pareto-based acquisition strategies for multi-objective Bayesian optimization (MOBO) in molecular design tasks. Rather than contrasting MOBO with scalarization broadly—which encompasses a diverse array of formulations including random scalarizations [golovin2020randomhypervolumescalarizationsprovable, paria2019flexibleframeworkmultiobjectivebayesian]—we focus on a controlled comparison between two representative acquisition strategies: Expected Hypervolume Improvement (EHVI), which explicitly targets Pareto front expansion, and a fixed-weight Expected Improvement (EI), a baseline scalarized acquisition function. We aim to answer the following questions:
@@ 3.3 Evaluation Metrics and Performance Indicators
To evaluate the performance of multi-objective optimization across both convergence quality and molecular diversity, we adopted three complementary metrics: Hypervolume Indicator (HVI) [fonsecahypervolume], the R2 indicator [hansenr2] and the #Circles metric [xie2023spaceexploredmeasuringchemical]. The HVI measures the volume of the objective space dominated by the non-dominated solutions relative to a reference point, capturing both convergence to the Pareto front and diversity of trade-offs explored. A higher HVI value indicates broader and more optimal coverage of the objective space. The $R^{2}$ indicator quantifies the quality of the Pareto front approximation by comparing it to a fixed set of uniformly distributed reference directions, as adapted from the setup in jain2023multi. Specifically, it computes the augmented Tchebycheff scalarization in each direction: for each reference vector $v$, it calculates the worst-case deviation $\text{max}_{i}v_{i}\cdot|u_{i}-s_{i}|$, where $u$ is the utopian (ideal) point and $s$ is a candidate solution. It then selects the best such solution for each direction and averages over all directions. Thus, lower $R^{2}$ values indicate a front that is both closer and more uniformly distributed relative to the ideal front [hansenr2]. Finally, the #Circles metric quantifies structural diversity in the chemical space by counting the number of pairwise dissimilar molecules exceeding a Tanimoto distance threshold $t$. We compute it on the set of Pareto-optimal candidates from the initial and acquired molecules over 200 BO evaluations, as the metric is designed to assess the effective space explored by a representative set of high-quality or relevant solutions rather than all generated samples [xie2023spaceexploredmeasuringchemical]. Informed by axiomatic diversity principles, #Circles offers a geometry-aware, thresholded view of chemical diversity: at higher thresholds (i.e. requiring greater dissimilarity), a larger #Circles value indicates that the candidate set spans more structurally distinct regions of the chemical space.
@@ 4 Results
This section presents a comparative evaluation of EHVI and scalarized EI across 3 multi-objective molecular optimization tasks. Across all tasks and metrics - hypervolume (Figure 1), $R^{2}$ indicator (Figure 2), and chemical diversity via the #Circles metric (Figure 3), EHVI demonstrates consistent advantages. Statistical tests provided in Appendix A.2 (Tables 3, 4) corroborate these trends: for hypervolume, EHVI achieves medium to large effect sizes (Cohen’s d=0.5761.093d=0.576-1.093) and favorable Cliff’s delta values indicating better performance across matched trials. For the $R^{2}$ indicator, lower values for EHVI suggest improved approximation of the Pareto front, with strong negative effect sizes (e.g. d=2.56d=-2.56 on Fexofenadine MPO). In terms of diversity, EHVI consistently explores more structurally distinct solutions at higher Tanimoto thresholds. These results demonstrate EHVI’s robustness: it not only accelerates convergence and enhances front coverage, it also maintains better front approximation and promotes chemically diverse candidate solutions. We next present the detailed metric-wise breakdown across all tasks.
@@ 5 Discussion
Our findings demonstrate that EHVI yields more sample-efficient exploration across a range of realistic multi-property optimization (MPO) tasks. These gains are reflected across three orthogonal metrics—hypervolume coverage, front approximation accuracy ($R^{2}$), and structural diversity—and are statistically robust across random seeds, as shown in Appendix A.2 (Tables 3, 4). Crucially, these improvements are not due to changes in surrogate fidelity, kernel choice, or representation capacity: both EHVI and scalarized EI operate under the same model assumptions and input features, underscoring the importance of acquisition strategy alone in driving performance.
@@ 6 Conclusion
We conducted a focused comparison between Pareto-based Bayesian optimization (EHVI) and scalarized Expected Improvement (EI) with fixed weights across multi-objective molecular design tasks. EHVI consistently outperforms EI, achieving higher hypervolume, lower $R^{2}$ values, and greater structural diversity. Statistical effect size analysis further confirms these gains are meaningful despite limited runs. Unlike scalarized EI, which optimizes a fixed utility, EHVI explicitly captures trade-offs, leading to more balanced and diverse Pareto fronts. These results demonstrate the value of Pareto-aware acquisition in data-constrained settings and motivate future work on adaptive surrogates, representation learning, and diversity-aware BO strategies.
Acknowledgements
This work was conducted as part of the first author’s Master’s thesis at University College London. We thank the thesis supervisors and committee members for their guidance and feedback.
Appendix A Appendix
A.1 Final Performance Metrics after 200 BO evaluations
@@
A.2 Cohen’s $d$ and Cliff’s Delta
@@
Fexofenadine 2.560-2.560 1.000-1.000
Amlodipine 0.770-0.770 0.556-0.556
Perindopril 1.602-1.602 0.778-0.778